OpenStudy (anonymous):

If f(x) = 3/5x and g(x) = - 3x/2 find f(x) + g(x)

OpenStudy (whpalmer4):

Is that \[f(x) = \frac{3}{5}x\] or \[f(x) = \frac{3}{5x}\]?

OpenStudy (anonymous):

The second one.

OpenStudy (whpalmer4):

\[\frac{3}{5x}+(-\frac{3x}{2})\]Okay, you need a common denominator. What will you use?